# suppSupplementary References

### v2

@@ 4.1 Properties of the Oracle Procedure
Proof. We first explain what the statement entails and then prove it. The statement says that decisions in quadrants I and III of Table 1 are exact-optimal. We demonstrate that starting from any decision not matching the claimed optimal decision in quadrants I and III and switching to the claimed optimal yields a lower FDR and a higher objective function. The decisions in quadrants II and IV are together rank-optimal. To prove that, among and across quadrants II and IV, we show that not altering or respecting the base decision rules by ranking, i.e., to exclude first in quadrant II and to include first in quadrant IV, leads to a suboptimal procedure. In other words, switching from any final decision that does not match our proposed procedure to one that matches it will lead to a lower FDR and a higher objective function value.
@@ 4.2 Second-Order Mean Correction
While we approximate $\ln\hat{p}_{1}/\hat{p}_{0}$ to a normal distribution as in Katz et al. (1978), we recommend doing a further second-order mean correction because our simulations show that a standard normal as derived by (Katz et al., 1978) can be made more accurate.
Let $10<p_{1}<$ 1, then define $f(p)=\ln p$, and by the second order approximation $f(p)\approx f(p_{1})+(p-p_{1})f^{\prime}(p_{1})+(p-p_{1})^{2}f^{\prime\prime}(p_{1})/2$. Plugging in $p=\hat{p}_{1}$, we get $\ln\hat{p}_{1}\approx\ln p_{1}+(\hat{p}_{1}-p_{1})f^{\prime}(p_{1})+(\hat{p}_{1}-p_{1})^{2}f^{\prime\prime}(p_{1})/2$. Where $f^{\prime\prime}(p_{1})=-1/p_{1}^{2}$. Thus the second order term is $-(1/2)(\hat{p}_{1}/p_{1}-1)^{2}$ which in expectation equals $(-1/2)Var(\hat{p}_{1}/p_{1})$. The result for $10<p_{0}<$ 1 is similar, i.e., $E\ln(\hat{p}_{0}/p_{0})\approx-(1/2)Var(\hat{p}_{0}/p_{0})$ and not zero as is usually assumed.
The correction procedure would add estimates of $(1/2)Var(\hat{p}_{1}/p_{1})$ and $(1/2)Var(\hat{p}_{0}/p_{0})$ when computing the lfdr. Our simulations found that this correction approximates the logarithm function better. [^3] A similar correction to the variance estimate can be applied. It will require further derivations but will likely not impact precision much.
@@ 4.3 Properties of the Data-Driven Procedure
In practice, the lfdr statistic has to be estimated using the data. There are a few approaches to estimating the lfdr statistic: by Efron et al. (2015) or Storey et al. (2024), Sun and Cai (2007) via Jin and Cai (2007), and Gaussian mixture models (see Fraley and Raftery (2002) for a general exposition). Because the lfdr is estimated, we prove that the data-driven approach asymptotically maintains the level of control of the FDR as the oracle approach. Later, we use numerical simulation to demonstrate the procedure’s finite sample performance.
@@
Proof. Define the estimated weight of hypotheses $i$ as $\widehat{W}_{i}:=(\widehat{lfdr}_{i}-\alpha)\cdot c_{i}$ and the ratio of value to weight $\widehat{R}_{i}:=\Pi_{i}\cdot\{\widehat{lfdr}(h_{i})/\widehat{lfdr}(h_{i}+\hat{\sigma}_{i})-1\}/\widehat{W}_{i}$. To simplify the proof, we assume w.l.o.g that $c_{i}\equiv 1$ and $\Pi_{i}\equiv 1$.
@@
By construction, the data-driven procedure ensures $\mathbb{E}_{\widehat{lfdr}}\{\sum_{i\in\mathcal{R}_{K}}(1-\theta_{i})/K\}\leq\alpha$ where $\mathcal{R}_{K}$ is the set of rejected hypotheses and $K$ is the number of rejections. This guarantee only holds for the estimated lfdrs, and now we show that it will hold for
@@
which holds because $|\widehat{ldfr}_{i}-lfdr_{i}|=o_{P}(1)$.444 This assumption is similar to those in the results of Basu et al. (2018), Appendix B.2.
@@ 5.1 Synthetic Data
Using simulation, we compare our procedure to the one-sided no correction and two FDR control procedures: adaptive BH (Benjamini et al., 2006), and SC (Sun and Cai, 2007). Both control procedures do not use the value of the lift in their objective. Because our method does take the lifts into account in the objective, we expect a significant increase in total lifts with the same level of the FDR when using our method compared to the other two. Berman and Van den Bulte (2022) inspired part of our setting. However, we vary the baseline (aka null) conversion rates and fix the difference from the non-null conversion rates (the effect size). Hence, experiments with lower baseline conversion rates generate larger lifts. All simulation results we report are an average of 400 replications. We assume each experiment is an A/B test with one baseline (control) and one treatment variation.
@@
Figure 2 provides a sample realization of the baseline conversion rates and profits. The setup with Gamma(1, 1) corresponds closely to the unweighted cases. We expect a higher objective function or total profit in the skewed cases from incorporating the weights. Because of the weighted nature of the heterogeneity, we further compare the performance of our procedure with the weighted FDR control procedure of Basu et al. (2018), which we denote BCDS. We expect our procedure to perform better because the total lifts are taken into account directly when controlling for the FDR.
@@
The figure displays the realized baseline conversion rates and profits of 2,000 experiments from the following Beta and Gamma distributions. [Upper panel, left to right] Baseline conversion rates are distributed Beta with parameters (1, 1), (1, 0.5), and (1, 0.25), respectively. Note that the top left plot corresponds to a realization from the uniform distribution on [0, 1]. Moving further to the right makes the distribution increasingly left-skewed. [Lower panel, left to right] Baseline profit distributions are Gamma with parameters (1, 1), (1/4, 1/4), and (1/9, 1/9), respectively. Note the standard deviation increases from one to three while keeping the mean profits at one for all the distributions.
@@
A higher percentage of power or profit for uncorrected tests is expected. In cases when it is hard to distinguish false from true discoveries, higher false rejections often mean more true rejections, leading to higher power. The argument is similar for the profit as true rejections will have a positive lift. We also report the net profit (i.e., profit after deducting the weighted costs) for the last simulation instance to give a sense of the economic benefit of incorporating FDR-based metrics. For illustrative purposes, we chose a mean cost-to-profit ratio of 2%. Although the method does not optimize the net profit directly, we see that RBL achieves a higher net profit than all other methods, improving the uncorrected method by more than 60% and by about 8-12% from BH and BCDS methods.
In the empirical application, we cannot observe the true lifts and hence cannot compute the profit. However, we can compute the observed average lift per rejection. Assuming the FDR methods we compare to control the FDR at the same level, a better algorithm will generate fewer rejections or reject hypotheses with higher lifts. Hence, although in the empirical application, we will use lift estimates, which are noisy measurements of the true lifts, a higher lift per rejection rate likely indicates that the algorithm will generate a higher profit. This pattern is evident in Table 2 where we observe that a higher net profit also implies a higher average lift per rejection.
@@ 5.2 Application: Optimizely Experiments
The data includes results from 2,766 experiments on the Optimizely platform starting in April 2014 (Berman and Van den Bulte, 2022). The experiments were all website experiments conducted by more than 1,000 experimenting companies. Every experiment had one control (baseline) condition and one or more variations of treatments. In total, there are results from 4,965 variations in our data. Because the lifts in experiments with many variations and one baseline are possibly correlated, we include non-baseline variations selected randomly with a probability proportional to the inverse of the number of variations in each experiment. This random selection may contain a mild dependency and gives us about 2,700 experiments to analyze. For each experiment, we have data about the total number of unique online visitors exposed to a specific variation of a webpage (corresponding to $n_{i0}$ and $n_{i1}$) and the total number of clicks on that page, which is the dependent variable in the experiment. These values allow us to compute the conversion rate (in terms of clicks) of the baseline control and the non-baseline variation. These conversion rates are used to compute lifts, z-scores, and p-values for the lifts.
Similar to our synthetic data analysis, we compare the RBL method with the Uncorrected, BH, SC, and BCDS methods. In this case, the ground truth (true population level conversion rates) is unknown, so we cannot report the FDR, power, or profit. We report the number of rejections for each method at different nominal level choices of 0.01, 0.05, and 0.10. In this dataset, there is no heterogeneity in baseline profit per conversion. Hence, we initially do analyses without heterogeneity in baseline profits but then extend them by generating heterogeneous profits from the gamma distribution as before.
@@
The table compares the performance of the RBL procedure to the Uncorrected, BH, SC, and BCDS methods. The scale parameter for the gamma distribution for the profits is varied in $\theta_{\gamma}\in\{\text{NA},1,2\}$, where ‘NA’ indicates the homogeneous baseline profit scenario. We implement the FDR control methods at a nominal $\alpha=\{0.01,0.05,0.10\}$. We set the random number generator seed for the selection of experiments and the profit weight generation at 2766.
@@
$\theta_{\gamma}=\text{NA}$ Rejections 179 111 111 112 105
@@
$\theta_{\gamma}=\text{NA}$ Rejections 324 142 145 145 133
@@
$\theta_{\gamma}=\text{NA}$ Rejections 475 169 171 172 153
@@
$\theta_{\gamma}=1$ Rejections 475 169 171 161 148
@@
$\theta_{\gamma}=1$ Rejections 324 142 145 134 125
@@
$\theta_{\gamma}=2$ Rejections 324 142 145 127 123
@@
In Table 3, we observe that the Uncorrected method rejects many hypotheses, with an average observed lift of approximately 33%. From our simulation study, we expect that 40%-50% of these rejections may be false positives. Practitioners often use an FDR control procedure such as the BH, SC, or BCDS methods to counter that. In both the simulation and application, we expect the performance of the BH and the SC methods to perform similarly, as they both guarantee FDR control at the chosen nominal level. Indeed, we observe this behavior in our application. The BCDS method is nearly identical in the homogeneous baseline profit scenario. In the heterogeneous (weighted) scenario, the BCDS method has higher weighted estimated lifts and higher average lift per rejection. Our proposed RBL method achieves the highest estimated lifts and average lift per rejection in all scenarios.
@@
6 Related Work
This section positions our work within the extant literature on A/B testing and FDR control. Azevedo et al. (2020) consider a framework to screen innovations when the distribution of the underlying effects can be fat-tailed. They do not consider false discoveries, the costs of implementation, or the heterogeneity in baselines that our method addresses. Azevedo et al. (2019) provide an overview of classical and recent methods available in this context. Their summary primarily discusses estimation methods leading up to and until nonparametric empirical Bayes estimation strategies. Finally, Gu and Koenker (2023) discusses the ranking and selection of a pre-determined number of top interventions in a decision-theoretic framework, while our work considers the implementation of all ‘profitable’ interventions, taking their costs into account.
Another consideration in the practical implementation of A/B is their online or sequential characteristics. Johari et al. (2022) implement and analyze elements of always valid p-values and confidence intervals under sequential testing. Zhou et al. (2023) discuss optimal boundaries for stopping rules. Chiong and Joo (2023) minimize the maximum regret specifically for a multi-arm online experiment. Feit and Berman (2019) analyze the choice of sample size that maximizes cumulative profits. Bartoš and Maier (2022) suggest decreasing the nominal level to reduce FDR, while Schultzberg et al. (2024) discuss several business metrics implemented at Spotify. Deng et al. (2016) advocates Bayesian testing to justify continuous monitoring. Gronau et al. (2019) also discuss Bayesian approaches to testing to incorporate prior expert beliefs on the competing hypotheses. These works focus on single test scenarios, not multiple or large-scale experimentation scenarios. We refer readers to Larsen et al. (2024) for an overview of statistical challenges in A/B testing.
@@ 7 Discussion
Future work can look into optimizing the experimental design while considering that an FDR procedure will be used during the analysis. For example, the number of variations and the sample size affect the cost of the experiment and the efficacy of the FDR control method. How to best incorporate these considerations in our statistical decision-making framework has promising potential. Another potential avenue of inquiry is to analyze issues arising in practical A/B testing due to optional stopping, peeking, or other post-hoc adjustments that occur when the tests are online. Finally, because it is recommended to perform replication following an FDR control procedure (see, for example, Berman and Van den Bulte (2022)), a future avenue for research is to understand how FDR control procedures and replications should be designed in conjunction to avoid the wastage of resources.

### v3

@@ 4.1 Properties of the Oracle Procedure
Proof. We first explain what the statement entails and then prove it. The statement says that decisions in quadrants I and III of Table 1 are exact-optimal. We demonstrate that starting from any decision not matching the claimed optimal decision in quadrants I and III and switching to the claimed optimal decision yields a lower FDR and a higher objective function. The decisions in quadrants II and IV are together rank-optimal. To prove that, among and across quadrants II and IV, we show that not altering or respecting the base decision rules by ranking, i.e., to exclude first in quadrant II and to include first in quadrant IV, leads to a suboptimal procedure. In other words, switching from any final decision that does not match our proposed procedure to one that matches it will lead to a lower FDR and a higher objective function value.
@@ 4.2 Second-Order Mean Correction
While we approximate $\ln\hat{p}_{1}/\hat{p}_{0}$ to a normal distribution as in Katz et al., (1978), we recommend doing a further second-order mean correction because our simulations show that a standard normal as derived by (Katz et al., , 1978) can be made more accurate.
Let $0<p_{1}<1$, then define $f(p)=\ln p$, and by the second order approximation $f(p)\approx f(p_{1})+(p-p_{1})f^{\prime}(p_{1})+(p-p_{1})^{2}f^{\prime\prime}(p_{1})/2$. Plugging in $p=\hat{p}_{1}$, we get $\ln\hat{p}_{1}\approx\ln p_{1}+(\hat{p}_{1}-p_{1})f^{\prime}(p_{1})+(\hat{p}_{1}-p_{1})^{2}f^{\prime\prime}(p_{1})/2$. Where $f^{\prime\prime}(p_{1})=-1/p_{1}^{2}$. Thus the second order term is $(1/2)(\hat{p}_{1}/p_{1}-1)^{2}$ which in expectation equals $(-1/2)Var(\hat{p}_{1}/p_{1})$. The result for $0<p_{0}<1$ is similar, i.e., $E\ln(\hat{p}_{0}/p_{0})\approx-(1/2)Var(\hat{p}_{0}/p_{0})$ and not zero as is usually assumed.
The correction procedure would add estimates of $(1/2)Var(\hat{p}_{1}/p_{1})$ and $(1/2)Var(\hat{p}_{0}/p_{0})$ when computing the lfdr. Our simulations found that this correction approximates the logarithm function better. [^4] A similar correction to the variance estimate can be applied. It will require further derivations, but will likely not impact precision much.
@@ 4.3 Properties of the Data-Driven Procedure
In practice, the lfdr statistic has to be estimated using the data. There are a few approaches to estimating the lfdr statistic: by Efron et al., (2015) or Storey et al., (2024), Sun and Cai, (2007) via Jin and Cai, (2007), and Gaussian mixture models (see Fraley and Raftery, (2002) for a general exposition). Because the lfdr is estimated, we prove that the data-driven approach asymptotically maintains the level of control of the FDR as the oracle approach. Later, we use numerical simulation to demonstrate the procedure’s finite sample performance.
@@
Proof. Define the estimated weight of hypotheses $i$ as $\widehat{W}_{i}:=(\widehat{lfdr}_{i}-\alpha)\cdot c_{i}$ and the ratio of value to weight $\widehat{R}_{i}:=\Pi_{i}\cdot\{\widehat{lfdr}(h_{i})/\widehat{lfdr}(h_{i}+\hat{\sigma}_{i})-1\}/\widehat{W}_{i}$. To simplify the proof, we assume w.l.o.g. that $c_{i}\equiv 1$ and $\Pi_{i}\equiv 1$.
@@
[…]
@@
which holds because $|\widehat{ldfr}_{i}-lfdr_{i}|=o_{P}(1)$.555 This assumption is similar to those in the results of Basu et al., (2018), Appendix B.2.
@@ 5.1 Synthetic Data
Using simulation, we compare our procedure to the one-sided no correction and two FDR control procedures: adaptive BH (Benjamini et al., , 2006), and SC (Sun and Cai, , 2007). Neither control procedures use the value of the lift in its objective. Because our method takes the lifts into account in the objective, we expect a significant increase in total lifts with the same level of the FDR when using our method compared to the other two. Berman and Van den Bulte, (2022) inspired part of our setting. However, we vary the baseline (aka null) conversion rates and fix the difference from the non-null conversion rates (the effect size). Hence, experiments with lower baseline conversion rates generate larger lifts. All simulation results we report are the average of 400 replications. We assume each experiment is an A/B test with one baseline (control) and one treatment variation.
@@
Figure 2 provides a sample realization of the baseline conversion rates and profits. The setup with Gamma(1, 1) corresponds closely to the unweighted cases. We expect a higher objective function or total profit in the skewed cases from incorporating the weights. Because of the weighted nature of the heterogeneity, we further compare the performance of our procedure with the weighted FDR control procedure of Basu et al., (2018), which we denote BCDS. We expect our procedure to perform better because the total lifts are taken into account directly when controlling for the FDR.
@@
The figure displays the realized baseline conversion rates and profits of 2,000 experiments from the following Beta and Gamma distributions. [Upper panel, left to right] Baseline conversion rates are distributed Beta with parameters (1, 1), (1, 0.5), and (1, 0.25), respectively. Note that the top left plot corresponds to a realization from the uniform distribution on [0, 1]. Moving further to the right makes the distribution increasingly left-skewed. [Lower panel, left to right] Baseline profit distributions are Gamma with parameters (1, 1), (1/4, 1/4), and (1/9, 1/9), respectively. Note that the standard deviation increases from one to three while keeping the mean profits at one for all the distributions.
@@
A higher percentage of power or profit for uncorrected tests is expected. In cases when it is hard to distinguish false from true discoveries, higher false rejections often mean more true rejections, leading to higher power. The argument is similar for the profit, as true rejections will have a positive lift. We also report the net profit (i.e., profit after deducting the weighted costs) for the last simulation instance to give a sense of the economic benefit of incorporating FDR-based metrics. For illustrative purposes, we chose a mean cost-to-profit ratio of 2%. Although the method does not optimize the net profit directly, we see that RBL achieves a higher net profit than all other methods, improving the uncorrected method by more than 60% and by about 8-12% from the BH and BCDS methods.
In the empirical application, we cannot observe the true lifts and hence cannot compute the profit. However, we can compute the observed average lift per rejection. Assuming the FDR methods we compare to control the FDR at the same level, a better algorithm will generate fewer rejections or reject hypotheses with higher lifts. Hence, although in the empirical application, we will use lift estimates, which are noisy measurements of the true lifts, a higher lift per rejection rate likely indicates that the algorithm will generate a higher profit. This pattern is evident in Table 2, where we observe that a higher net profit also implies a higher average lift per rejection.
@@ 5.2 Application: Optimizely Experiments
The data includes results from 2,766 experiments on the Optimizely platform starting in April 2014 (Berman and Van den Bulte, , 2022). The experiments were all website experiments conducted by more than 1,000 experimenting companies. Every experiment had one control (baseline) condition and one or more variations of treatments. In total, there are results from 4,965 variations in our data. Because the lifts in experiments with many variations and one baseline are possibly correlated, we include non-baseline variations selected randomly with a probability proportional to the inverse of the number of variations in each experiment. This random selection may contain a mild dependency and gives us about 2,700 experiments to analyze. For each experiment, we have data about the total number of unique online visitors exposed to a specific variation of a webpage (corresponding to $n_{i0}$ and $n_{i1}$) and the total number of clicks on that page, which is the dependent variable in the experiment. These values allow us to compute the conversion rate (in terms of clicks) of the baseline control and the non-baseline variation. These conversion rates are used to compute lifts, z-scores, and p-values for the lifts.
Similar to our synthetic data analysis, we compare the RBL method with the Uncorrected, BH, SC, and BCDS methods. In this case, the ground truth (true population level conversion rates) is unknown, so we cannot report the FDR, power, or profit. We report the number of rejections for each method at different nominal level choices of 0.01, 0.05, and 0.10. In this dataset, there is no heterogeneity in baseline profit per conversion. Hence, we initially do analyses without heterogeneity in baseline profits, but then extend them by generating heterogeneous profits from the gamma distribution as before.
@@
$\theta_{\gamma}=\text{NA}$ # Rejections 179 111 111 112 105
@@
$\theta_{\gamma}=\text{NA}$ # Rejections 324 142 145 145 133
@@
$\theta_{\gamma}=\text{NA}$ # Rejections 475 169 171 172 153
@@
$\theta_{\gamma}=1$ # Rejections 475 169 171 161 148
@@
$\theta_{\gamma}=1$ # Rejections 324 142 145 134 125
@@
$\theta_{\gamma}=2$ # Rejections 324 142 145 127 123
@@
The table compares the performance of the RBL procedure to the Uncorrected, BH, SC, and BCDS methods. The scale parameter for the gamma distribution for the profits is varied in $\theta_{\gamma}\in\{\text{NA},1,2\}$, where ‘NA’ indicates the homogeneous baseline profit scenario. We implement the FDR control methods at a nominal $\alpha=\{0.01,0.05,0.10\}$. We set the random number generator seed for the selection of experiments and the profit weight generation to 2766.
In Table 3, we observe that the Uncorrected method rejects many hypotheses, with an average observed lift of approximately 33%. From our simulation study, we expect that 40%-50% of these rejections may be false positives. Practitioners often use an FDR control procedure such as the BH, SC, or BCDS methods to counter that. In both the simulation and application, we expect the performance of the BH and the SC methods to perform similarly, as they both guarantee FDR control at the chosen nominal level. Indeed, we observe this behavior in our application. The BCDS method is nearly identical in the homogeneous baseline profit scenario. In the heterogeneous (weighted) scenario, the BCDS method has higher weighted estimated lifts and a higher average lift per rejection. Our proposed RBL method achieves the highest estimated lifts and average lift per rejection in all scenarios.
@@
6 Additional Related Work
This section positions our work within the extant literature on A/B testing and FDR control. Azevedo et al., (2020) considers a framework to screen innovations when the distribution of the underlying effects can be fat-tailed. They do not consider false discoveries, the costs of implementation, or the heterogeneity in baselines that our method addresses. Azevedo et al., (2019) provides an overview of classical and recent methods available in this context. Their summary primarily discusses estimation methods leading up to and including nonparametric empirical Bayes estimation strategies. Finally, Gu and Koenker, (2023) discusses the ranking and selection of a pre-determined number of top interventions in a decision-theoretic framework, while our work considers the implementation of all ‘profitable’ interventions, taking their costs into account.
Another consideration in the practical implementation of A/B is their online or sequential characteristics. Johari et al., (2022) implement and analyze elements of always valid p-values and confidence intervals under sequential testing. Zhou et al., (2023) discusses optimal boundaries for stopping rules. Chiong and Joo, (2023) minimize the maximum regret specifically for a multi-arm online experiment. Feit and Berman, (2019) analyzes the choice of sample size that maximizes cumulative profits. Bartoš and Maier, (2022) suggest decreasing the nominal level to reduce FDR, while Schultzberg et al., (2024) discuss several business metrics implemented at Spotify. Deng et al., (2016) advocates Bayesian testing to justify continuous monitoring. Gronau et al., (2019) also discusses Bayesian approaches to testing to incorporate prior expert beliefs on the competing hypotheses. These works focus on single test scenarios, not multiple or large-scale experimentation scenarios. We refer readers to Larsen et al., (2024) for an overview of statistical challenges in A/B testing.
@@ 7 Discussion
Future work can look into optimizing the experimental design while considering that an FDR procedure will be used during the analysis. For example, the number of variations and the sample size affect the cost of the experiment and the efficacy of the FDR control method. How to best incorporate these considerations in our statistical decision-making framework has promising potential.
As mentioned in the introduction, our work focuses on experimentation programs (Sudijono et al., , 2024); however, another potential avenue of inquiry is to analyze issues arising in online A/B testing due to optional stopping, peeking, or other post-hoc adjustments. Finally, because it is recommended to perform replication following an FDR control procedure (see, for example, Berman and Van den Bulte, (2022)), a future avenue for research is to understand how FDR control procedures and replications should be designed in conjunction to avoid the wastage of resources.
@@
Appendix A Analysis of Alternate Objective Formulation
Recall that we aimed to solve:
$\max_{\delta_{i}\in\{0,1\}}\mathbb{E}\sum_{i}\Pi_{i}\ell_{i}\delta_{i}\text{~~subject to~~}\mathbb{E}\{\sum(1-\theta_{i})\delta_{i}c_{i}/\sum\delta_{i}c_{i}\}\leq\alpha.$
A perhaps more straightforward formulation is to directly maximize the expected net profit:
$\max_{\delta_{i}\in\{0,1\}}\mathbb{E}\sum_{i}\{\Pi_{i}\ell_{i}-c_{i}\}\delta_{i}.$
This section discusses the statistical challenges of power and the validity of doing so.
Even though this formulation is a compound decision problem, the oracle decisions are individual in nature owing to the separability of the objective function and independence of tests. The Bayes optimal solution in this setting is $\delta_{i}=1$ if and only if $c_{i}/\Pi_{i}<E(\ell_{i}|\text{data})$. We discuss the merits of our formulation over this decision rule through a stylized simulation, with settings that closely match experimentation programs (Sudijono et al., , 2024) and our Optimizely data.
Figure 3 illustrates the results. For a set seed (1), we generate 2000 samples of the log risk-ratio with 80% coming from the null of $N(0,\cdot)$ and 20% from a non-null of $N(0.1,\cdot)$. The variance is chosen to match the log risk-ratio asymptotic, assuming 5000 visitors and a conversion rate of 10% for experiment versions A and B each. The cost-to-profit ratio takes the values of 2% or 10%, with the profit being fixed at five. Two oracle methods are compared, with our method being invariant to the scale of costs, and the Bayes optimal method for both values of the proportional cost are implemented. (Note that in the main text, all results in Section 5 are for the data-driven methodologies. Only this section compares the oracle methods and uses the same lfdr test statistics for both procedures for conceptual exploration.)
We chose a nominal FDR level for our method of 0.1. Figure 3 shows that the false discovery proportion is roughly maintained, and our method makes roughly 80 rejections, and about 70 of them are correct decisions. In contrast, we see that, depending on the cost-to-profit ratio, the Bayes optimal method makes either too many rejections, resulting in a large proportion (more than half) of them being false, or it rejects too few hypotheses, thereby also making the effort of running so many tests quite wasteful. [^7] We report the results for one seed value: they do not vary much for arbitrary choices of seeds. Overall, if experimenters are interested in long-term learning from experimentation, simply optimizing the short-term net profit is not the best philosophy to rely on, as it can be very wasteful or too erroneous. In contrast, if the implementation of the challenger intervention is viewed as an investment, then controlling for the ‘wasted’ proportion of total investment could be a reasonable constraint while maximizing the profit.
Our empirical observation is similar to that in Figure 7a of (Sudijono et al., , 2024) where the authors identify a range of p-value thresholds from liberal to conservative that are equivalent to the Bayes optimal decision rule, when compared to the nominal level, as the cost of implementation increases. We further emphasize that while calculating the cost of implementation is relatively straightforward, costs associated with human labor, as noted by the authors, costs associated with loss of correct learning, or cumulative costs associated with wrong knowledge are all expensive and uncertain for the firm. These costs cannot be easily measured, thereby leading to the additional difficulty of including them in a simplified objective function as discussed in this section. The idea of controlling the weighted FDR thus becomes more relevant owing to the more subtle interpretation of costs and false discoveries.
Figure 3: Illustration of the rejection sets for our and alternate methods.
The figure displays the rejection set for our method when compared to the rejection sets of the Bayes optimal decision rule. The log risk-ratio process is assumed to be from a normal mixture with a 0.8 probability of being the null of $N(0,\cdot)$ and the non-null distribution is $N(0.1,\cdot)$. The variances are appropriately chosen. We choose $\alpha=0.1$ for finding the rejection sets. We note that when the implementation cost-to-profit is a realistic 2%, the Bayes optimal decision rule makes many more null rejections than is desirable. Conversely, when the implementation cost-to-profit ratio is higher, the Bayes optimal procedure makes hardly any rejections, making the A/B testing effort wasteful.